OpenStudy (anonymous):

Hey! So im learning Inverse of trig functions in school, so can anyone please explain them to me? I was absent in class and I have no idea on how to even attempt the problem, heres how to do it but can someone explain why? Thank you in advance! :) Sinx + 1 = 0

OpenStudy (amistre64):

do you know what an inverse function is spose to do to begin with?

OpenStudy (anonymous):

yes :)

OpenStudy (amistre64):

good, inverse trig functions are the result of restricting the domain of thier normal trig function

OpenStudy (amistre64):

Sinx + 1 = 0 first thing you would want to do is get sinx all by itself; how would we do that?

OpenStudy (anonymous):

bring it to the other side so sinx = -1

OpenStudy (amistre64):

correct, and all that is left is to inverse the trig function, on both sides, to get the x out

OpenStudy (amistre64):

arcsin(sinx) = arcsin(-1) x = arcsin(-1)

OpenStudy (anonymous):

thats where i am confused! :/

OpenStudy (anonymous):

can you explain that?

OpenStudy (ghazi):

let's say there is a function y=sin x , then expressing y in terms of x means inverse of it , which is\[x= \sin ^{-1} y\]

OpenStudy (amistre64):

trig functions take angles as inputs and give ratios as outputs; the inverse trig functions therefore take in ratios and give out angles.

OpenStudy (amistre64):

if you know what an inverse function is, then there isnt much mystery to trig inverses.

OpenStudy (anonymous):

but the attachment i put above is the answer

OpenStudy (amistre64):

sinx=-1 is asking, what angles give us a sin ratio of -1 x = arcsin(-1) is asking or telling us the same thing; x equates to all the angles that have a sin ratio of -1

OpenStudy (amistre64):

youre most likely confused at why the answer is given in terms of more than one angle then?

OpenStudy (anonymous):

how did they get 3/2 ?

OpenStudy (amistre64):

can you give me one single angle that has a sin ratio of -1?

OpenStudy (amistre64):

3pi/2 is a radian measure that corresponds to 270 degrees

OpenStudy (amistre64):

radians are the prefered means of working with trig functions

OpenStudy (amistre64):

2pi = 360 degrees 270 is 3/4 of 360 degrees 2pi * 3/4 = 3pi/2

OpenStudy (ghazi):

its \[\frac{ 3 \pi }{ 2 }\] sorry amistre to interrupt you :)

OpenStudy (amistre64):

no problem, by all means i usually need interupting just to keep meself inline ;)

OpenStudy (ghazi):

:)

OpenStudy (amistre64):

since 3pi/2 is one angle that has a sin ratio of -1; and all angles that periods of 2pi from it have the same sin ratio; we can include all of the angles by adding on that little 2pi k part. where k is any integer

OpenStudy (amistre64):

it like saying the x^2 = 4 has the solutions +- 2

OpenStudy (anonymous):

but how is 3pi/2 an angle that has a a sin ratio of -1?

OpenStudy (amistre64):

what is the sin of 270?

OpenStudy (amistre64):

270 degrees that is

OpenStudy (amistre64):